At his utility maximizing bundle, Jim has marginal utility of 10 for beef snacks and marginal utility of 30 for carrot juice. Which of the following is true?

Question 32 options:

1)

Jim receives higher total utility from carrot juice.

2)

Jim receives higher total utility from beef snacks.

3)

Jim is willing to trade 3 units of carrot for one more beef snack.

4)

Jim’s MRS of carrot juice for beef snacks is 300

5)

Carrot juice costs three times as much as beef snacks.
Answer #1

Option 5 is correct answer

Marginal rate of substitution between beef snacks and carrot juice is the number of carrot juice given up for the additional unit of beef snack

therefore MRS=MU(beef stock)/MU(carrot juice)=10/30=1/3

Hence it means that we have to forgo 1/3rd of Carrot juice for additional unit of beef snack to maintain the same utility

Hence option 3 is not correct

At equilibrium MRS= Price ratio=1/3

Beef Snack Price/Carrot Juice Price=1/3

Carrot Juice Price=3* Beef Snack Price

Hence option 5) is correct